The displacement of a body is given by s=1/2gt2 where g is acceleration due to gravity the velocity of the body at an any time t is

Grade:12th pass

3 Answers

Vikas TU
7 years ago
velocity of the body at any time t would be:
differentiation of s  w.r.t  t.
that is:
ds/dt = gt
v = gt m/s.
Mukesh
6 years ago
The displacement of a body is given by 2s=gt×t where g is constant. The velocity of the body at any time t is
Vanshika
5 years ago
S= 1/2 gt square.
 So, U = zero.
 We know that V = U + at 
                                 U= zero
                           V= at . Here a = gravity (g)
                                        So, v = gt